OpenStudy (nilankshi):

cos A-sin A+sin^2 +1/cos A+sin A-1=cosec A+cot A

OpenStudy (nilankshi):

plzz...help mee..

OpenStudy (lgbasallote):

prove?

OpenStudy (nilankshi):

yes

OpenStudy (mimi_x3):

\[\huge \frac{cosA-sinA+\sin^{2}+1}{cosA+sinA-1} = cosecA + cotA\] ? does not look right..

OpenStudy (anonymous):

Is everything before the "/" on the top and everything after on the bottom?

OpenStudy (nilankshi):

noo..its sin A +1 not sin ^2

OpenStudy (mimi_x3):

\[\large \frac{cosA-sinA+1}{cosA+sinA-1} \] ?

OpenStudy (nilankshi):

yes..

OpenStudy (nilankshi):

plzz..hep mee..

OpenStudy (mimi_x3):

Try expanding the RHS first. \[cosecA = \frac{1}{sinA} \] \[cotA = \frac{cosA}{sinA} \]

OpenStudy (nilankshi):

hmm..

OpenStudy (nilankshi):

ten..what we can doo..??

OpenStudy (nilankshi):

then

OpenStudy (apoorvk):

first please confirm which question is right? is what mimi wrote above correct?

OpenStudy (nilankshi):

yes..